OpenStudy (anonymous):

Can i get atleast one physical situtation which is depicted and interpreted by differential equation

OpenStudy (wrwrwr):

A few examples from the course: Newton's law of cooling: \[T' = k (T_e - T)\] where T is temperature of some object, T_e is external temperature and k a coupling constant (conductiveness / isolation). Newton's law of universal gravitation: \[r'' = G M / r^2\] where r is distance to the other mass (so r'' is acceleration), G is the gravitational constant and M is the other mass. Mass attached to a spring with some damping (e.g. Hooke's law): \[m x'' + b x' + k x = f(t)\] where m is the mass, x is its position, b is a damping constant, k is spring's strength and f is some driving force. Circuit with resistor, inductor and capacitor (RLC): \[L Q'' + R Q' + Q / C = V\] where Q is charge on the capacitor, L is inductance, R resistance, C capacitance and V a driving voltage.

OpenStudy (anonymous):

thank u .......